a researcher collecting data on college students finds that
21 have a printer
56 have a computer and
12 have both
a) Why might a student have a printer but not have a computer?
b) If a student has a computer what is the probability that the student also has
a printer

Answer by Edwin McCravy(20054)

a researcher collecting data on college students finds that
21 have a printer
56 have a computer and
12 have both
a) Why might a student have a printer but not have a computer?
Because with a printer he/she can scan a page and then print it even 
if he/she has no computer.

b) If a student has a computer what is the probability that the student
also has a printer.
Suppose there are N students at the college.  The conditional probability
formula gives this:

P(has printer | has computer)  = P(has printer AND has computer)/P(has computer)

= (12/N)/(56/N) = (12/N)(N/56) = 12/56 = 3/14

Edwn